# Interplay between Majorana and Shiba states in a minimal Kitaev chain coupled to a superconductor

### v1

@@ I Introduction
The Kitaev chain [1] is the minimal model for one-dimensional spinless p-wave superconductivity. In the topological phase, the chain hosts Majorana Zero Modes (MZMs) [2, 3, 4, 5, 6] at its ends, which feature non-abelian statistics that can be exploited for topological quantum information processing [7, 8, 9, 10]. New platforms based on arrays of quantum dots (QDs) connected through mesoscopic superconductors [11, 12, 13, 14] have emerged as a promising route toward bottom-up engineering of artificial Kitaev chains. This approach largely avoids undesired effects associated to disorder and material inhomogeneities inherent to other platforms [15, 16, 17].
@@
A minimal Kitaev chain can be realized by just two QDs coupled through a mesoscopic superconductor that mediates crossed Andreev reflection (CAR) and elastic cotunneling (ECT) between the QDs [11, 12]. Such minimal Kitaev chains host “poor man’s Majorana states”, that share properties with their topological counterparts [18], but lack topological protection, since they appear at fine-tuned “sweet spots” in parameter space with equal CAR and ECT. Recent experiments have demonstrated exquisite control on CAR and ECT amplitudes [19, 20, 21, 22], which allows to tune sweet spots [23, 24, 22, 25, 26] where one MZM localizes in each QD of the minimal chain. Interestingly, longer chains, already at the three QD level, start to show a bulk gap and some degree of protection [26].
@@
So far, the assessment of sweet spots has been done via local and non-local normal conductance, by coupling the system to metallic leads [19, 20, 21, 24, 22, 25, 26]. In this work, we explore what happens beyond this minimal setup and theoretically study the transport properties of a minimal Kitaev chain coupled to a normal and a superconducting (SC) probe, see Fig.1(a), inspired by recent experimental configurations, Refs. 22, 26.
@@
Figure 1: (a) Sketch of a minimal Kitaev chain contacted to a normal (N) and a superconducting (S) lead, with tunneling strengths $\lambda_{n}$ and $\lambda_{s}$ respectively. (b) MZMs localization in the two QDs (blue and green) for different parameter regimes, assuming $t=\Delta_{p}$. (c-f) Normal conductance as a function of the QDs levels. Upper (Lower) panel shows the normal conductance as a function of $\varepsilon_{1}$ ($\varepsilon_{2}$). (d, f) Conductance cuts for different values of the quantum dots levels $\varepsilon/t=0,0.5,1,2$. We set $\Gamma_{n}/t=0.4$ and $\Gamma_{s}/t=0.1$ in all panels.
@@
One consequence of the replacement of one of the normal leads by a SC one, is the change of the the well-known quantized zero bias peak $G_{0}=2e^{2}/h$ [27, 28, 29], which now becomes $G_{s}(e|V|=\Delta)=(4-\pi)G_{0}$ when the BCS coherence peak aligns with the MZM [30, 31]. Despite the lack of topological protection in this minimal setup, we find a less obvious result: the universal conductance values through either the normal or the SC probe are robust against local detuning of the QD in contact with such probe. Further analysis in this work includes a detailed study of the role of the asymmetries and non-local effects induced by the inclusion of the non-trivial self-energy associated to the SC lead. As we discuss, this phenomenology could be used as a way to measure the degree of localization [32] of the MZMs in the minimal Kitaev chain and to characterize sweet spots.
@@
Interestingly, the regime of strong hybridization between the SC and the minimal Kitaev chain is governed by the emergence of Yu-Shiba-Rusinov (YSR) states [33, 34, 35, 36, 37, 38]. The interplay between MZMs and YSR has already been studied previously using minimal models neglecting the continuum of quasiparticles above the SC lead [39, 40]. Their inclusion, however, results in important physical consequences since the energy of the YSR states decreases when increasing the coupling to the lead, thus reducing the effective gap between MZMs and the excited states. A finite spin-polarization in the dots generates a destructive interference between the YSR states and the MZMs which spoils the topological properties of the latter, including the conductance quantization. In the language of non-Hermitian open quantum systems, we understand the topological properties of the hybrid junction in terms of Exceptional Point (EP) bifurcations in the subgap spectrum [41, 42, 43, 44, 45], characterized by a change of the MZMs localization.
@@
We finally offer, in Sec. VII, some conclusions summarizing the main results. Technical details like the explicit analysis of the poles of the system including the exceptional points, some notes on the Green functions used to compute both, spectral and transport properties and, the extension of the study of the physics associated to the QDs chemical potential to large couplings, are included in the appendices.
@@ II Modelization
We examine a minimal Kitaev chain coupled to a normal and a SC leads, as sketched in Fig.1(a). We consider that the system is subject to a strong magnetic field that fully polarizes the quantum dots spins. The two QDs are coupled via a central SC, that mediates ECT and CAR, with effective amplitudes $t$ and $\Delta_{p}$ respectively, realizing a one dimensional spinless p-wave SC that can host MZMs [11]. External gates can be used to tune the chemical potential on the two QDs independently ($\varepsilon_{1}$ and $\varepsilon_{2}$), changing the degree of localization of the MZMs, see Fig.1(b). The minimal Kitaev chain is coupled to the SC (normal) lead through the tunneling rate $\Gamma_{s}=\lambda_{s}^{2}/t_{s}$ ($\Gamma_{n}=\lambda_{n}^{2}/t_{n}$), being $\lambda_{s}$ ($\lambda_{n}$) the coupling and $t_{s}$ ($t_{n}$) the bandwidth.
@@
To describe the transport properties of the device, we use the non-equilibrium Green’s function formalism [46, 31, 47, 48, 49], that allows us to consider arbitrary coupling strengths between the minimal Kitaev chain and the leads, see Appendix D and F for details. In addition, the poles of the Green’s function determine the energy of the system’s states that can be related to the main transport features. Notice that for a spin-polarized chain, there is no subgap Andreev reflections with the SC probe [50] due to the incompatibility between order parameters across the junction [51, 30, 31] (i.e., a BCS lead with s-wave pairing $\Delta_{s}$ and a spinless SC with p-wave pairing $\Delta_{p}$ [52]).
In the sweet spot (i.e., $t=\Delta_{p}$ and $\varepsilon_{1}=\varepsilon_{2}=0$) and $\Gamma_{n}\rightarrow 0$ limit, the poles of the Green function are contained in the characteristic polynomial
$P(\omega)$ $=$ $\omega^{4}\big{[}\big{(}\Gamma_{s}^{2}\omega^{2}+(\Delta_{s}^{2}-\omega^{2})(\omega^{2}-4\Delta_{p}^{2})\big{)}^{2}$ (1)
$-4\Gamma_{s}^{2}(\Delta_{s}^{2}-\omega^{2})(\omega^{2}-2\Delta_{p}^{2})^{2}\big{]}.$
Along with the zero-energy poles, there appears other subgap structure that may describe the excited states (±$\pm z_{s,e}$) in certain parameter regimes. The non-perturbative coupling to the SC lead renormalizes the energy of such excited states, appearing as subgap poles detaching from the continuum. These poles can be obtained analytically (cf. App. A for details), and we reproduce here an approximation to the lowest contribution in $\Gamma_{s}$
$z_{s,e}\approx\frac{1}{\sqrt{3}}\sqrt{8\Delta_{p}^{2}+\Delta_{s}^{2}-\frac{\beta^{2}}{\alpha^{1/3}(\Gamma_{s})}-\alpha^{1/3}(\Gamma_{s})},$ (2)
where $\alpha(\Gamma_{s})=\big{(}\beta^{3}+12\sqrt{3}\,\Gamma_{s}\Delta_{p}^{2}\,\beta^{3/2}\big{)}$, and $\beta=(4\Delta_{p}^{2}-\Delta_{s}^{2})$.
@@
Figure 2: Local (a,b) and non-local (c,d) effects in SC transport in the weak coupling limit, varying the chemical potential of the dot in contact with the respective probe. First (Second) row shows the dependence of the SC conductance for various values of $\varepsilon_{1}$ ($\varepsilon_{2}$). The different columns show different coupling to the normal lead $\lambda_{n}/t=0.1$ (a,c) and $\lambda_{n}/t=0.5$ (b,d). Red dashed line marks the universal height value $G_{s}/G_{0}=(4-\pi)$. We set $\lambda_{s}/t=0.5$ in all panels.
@@ III Weak coupling
Figure 1(b) illustrates the Majorana localization when detuning the two QDs away from the sweet spot. When QD1 is detuned, the two MZMs have a finite spectral weight in QD2. Thus, when considering $\varepsilon_{1}\neq 0$, the normal lead couples to both MZMs, producing a destructive interference sharp dip at zero energy, Fig.1(c), as previously observed in Ref. 11. Moreover, due to electrons tunneling between the normal and the SC lead, mediated by the Kitaev chain, a peak appears at $e|V|=\Delta$, where $\Delta_{p}=\Delta_{s}=\Delta$, see also Fig.1(d).
By contrast, when QD2 is detuned from the sweet spot, the two MZMs acquire a finite weight in QD1. This leads to a quantized zero-bias conductance peak, showing a width that depends on $\varepsilon_{2}$, cf. Fig.1(f) and Ref. 11. Therefore, detuning either the QD attached to the normal lead or the one attached to the SC lead has very different effects on the transport properties.
Figure 4: (a) QD2 LDOS as a function of the hybridization parameter to the SC lead. (b,e) LDOS across the junction, showing the spreading of the states into the SC lead in the $\lambda_{n}/t\rightarrow 0$ limit. (c,f) LDOS cuts following the YSR state and the MZM respectively. (b,c) [(e,f)] Weak (Strong) coupling for $\lambda_{s}/t=0.5$ ($\lambda_{s}/t=4$) marked by arrows in (a). (d, g) Illustration of the weak and strong coupling across the junction (d) Weak coupling, minimal Kitaev chain. (g) Strong coupling, YSR state coupled to QD2 forming bonding and anti-bonding subgap states. We set $\lambda_{n}/t=1$ in the rest of the panels. White dashed line in panel (a) represents the analytic pole obtained from Eq. (1) when considering $\Gamma_{n}\neq 0$. The LDOS at QD1 and QD2 in (c,f) is rescaled for visualization purposes.
@@
Figure 2(c,d) illustrates the opposite behaviour, namely how SC transport is modified non-locally when the QD far from the SC lead is detuned. In this case, the lead couples to both MZMs. The resulting hybridization becomes evident when testing transport through the SC probe, where the quantized conductance value gets reduced as soon as $\varepsilon_{2}\neq 0$. Moreover, the coherent coupling between MZMs mediated by the SC lead, gives rise to subgap signatures in transport, Fig. 2(c). We remark that the subgap structure is not associated to any renormalization of subgap poles but to inelastic single-particle processes in transport, cf. Refs. [53] and App. D. When the coupling to the normal lead is increased, the spreading of the MZM into the normal lead enhances the SC transport, Fig. 2(d). Notice that this dependence with $\Gamma_{n}$ is the opposite to the one found in the $\varepsilon_{1}\neq 0$ case in Fig. 2(a, b).
@@ IV Strong coupling
Now we examine the electron transport as a function of the coupling to the SC lead, going beyond the weak coupling regime ($\Gamma_{s}/t\gtrsim 1$) where the excited state at $\omega=2\Delta$ is strongly renormalized by the lead, emerging as a subgap YSR state [53, 54, 55, 56, 39, 40]. Transport provides an ideal method to observe the interplay between the aforementioned YSR states and MZMs, Fig. 3. The onset of the YSR state appears as a new peak detaching from the continuum with linear dispersion around $\Gamma_{s}/t\gtrsim 0.5$, Figs. 3(a, c). This phenomenology does not affect the MZM signatures at $V\ll\Delta$, except for $\Gamma_{s}/t\gg 1$ and finite spin-polarization in QD1, see below and App. C. The conductance of the excited state is also quantized for $\Gamma_{s}/t\geq 1$ and $\varepsilon_{2}=0$. Intuitively, the excited state is no longer broadened by the SC lead when is well detached from the continuum, thus recovering the Majorana properties of the original excited state.
Detuning $\varepsilon_{1}$ barely affects the excited state and only leads to a dip in the zero-bias conductance, as commented in the weak coupling regime, see Fig. 1. The energy of the excited state is renormalized when $\varepsilon_{2}$ shifts away from the sweet spot, losing any quantized behaviour in transport. In particular, it gets pinned to $z_{s,e}\rightarrow\varepsilon_{2}$ for large $\Gamma_{s}$ values, cf. App. A.
@@
Figure 5: (a) QD2 LDOS as a function of the hybridization parameter to the SC lead increasing from weak to the ultra-strong coupling limit. First row, LDOS across the junction, showing the spreading of the states into the SC lead. Second row, real and imaginary part of the poles of the system (straight and dotted lines respectively) for $\lambda_{s}/t=10,60,100$ marked by arrows in (a). We set $\lambda_{n}/t=1$ in all panels. White dashed line in panel (a) represents the analytic pole obtained from Eq. (1) when considering $\Gamma_{n}\neq 0$.
@@
Transport through the SC lead is also affected when $\Gamma_{s}$ increases, Fig. 3(b, d). First, the universal height value $(4-\pi)$ moves to larger values of the voltage bias as the coupling to the SC lead is increased [53], illustrated by the yellow line. Figure 3(d) shows the onset of a subgap structure associated to the hybridization between Majoranas when detuning QD2 even in the weak coupling regime $\Gamma_{s}\rightarrow 0$, see also Fig. 2(c). This subgap structure is obtained from the single-particle contribution when testing transport through the SC probe when enabling thermally activated inelastic cotunneling, Thus, the hybridized MZMs behaves as a trivial zero energy subgap state in the strong tunneling regime [53], cf. App. D and E. Notice that the conductance peaks have a reduced value, compared to panel (b) and for $\Gamma_{s}/t\leq 1$ the dispersion of the subgap peak is linear with $\lambda_{s}$ (i.e., quadratic with $\Gamma_{s}$) highlighting the different origin with the observed subgap structure in Fig. 3(a,c).

### v3

@@ I Introduction
The Kitaev chain [1] is the minimal model for one-dimensional spinless p-wave superconductivity. In the topological phase, the chain hosts Majorana Zero Modes (MZMs) [2, 3, 4, 5, 6, 7] at its ends, which feature non-abelian statistics that can be exploited for topological quantum information processing [8, 9, 10, 11]. New platforms based on arrays of quantum dots (QDs) connected through mesoscopic superconductors [12, 13, 14, 15] have emerged as a promising route toward bottom-up engineering of artificial Kitaev chains. This approach largely avoids undesired effects associated to disorder and material inhomogeneities inherent to other platforms [16, 17, 18, 19].
@@
A minimal Kitaev chain can be realized by just two QDs coupled through a mesoscopic superconductor that mediates crossed Andreev reflection (CAR) and elastic cotunneling (ECT) between the QDs [12, 13]. Such minimal Kitaev chains host “poor man’s Majorana states”, that share properties with their topological counterparts [20], but lack topological protection, since they appear at fine-tuned “sweet spots” in parameter space with equal CAR and ECT. Recent experiments have demonstrated exquisite control on CAR and ECT amplitudes [21, 22, 23, 24], which allows to tune sweet spots [25, 26, 24, 27, 28] where one MZM localizes in each QD of the minimal chain. Interestingly, longer chains, already at the three QD level, start to show a bulk gap and some degree of protection [28, 29].
@@
So far, the assessment of sweet spots has been done via local and non-local normal conductance, by coupling the system to metallic leads [21, 22, 23, 25, 26, 24, 27, 28]. In this work, we explore what happens beyond this minimal setup and theoretically study the transport properties of a minimal Kitaev chain coupled to a normal and a superconducting (SC) probe, see Fig.1(a), inspired by recent experimental configurations, Refs. [24, 28].
@@
Figure 1: (a) Sketch of a minimal Kitaev chain contacted to a normal (N) and a superconducting (S) lead, with tunneling strengths $\lambda_{n}$ and $\lambda_{s}$ respectively. (b) MZMs localization in the two QDs (blue and green) for different parameter regimes, assuming $t=\Delta_{p}$. (c-f) Normal conductance as a function of the QDs levels. Upper (Lower) panel shows the normal conductance as a function of $\varepsilon_{1}$ ($\varepsilon_{2}$). (d, f) Conductance cuts for different values of the quantum dots levels $\varepsilon/t=0,0.5,1,2$. Inset shows a diagram of the setup, where the blue arrow marks the direction of the probing. We set $\Gamma_{n}/t=0.4$ and $\Gamma_{s}/t=0.1$ in all panels.
@@
One consequence of the replacement of one of the normal leads by a SC one, is the change of the well-known quantized zero-bias peak $G_{0}=2e^{2}/h$ [30, 31, 32], which now becomes $G_{s}(e|V|=\Delta)=(4-\pi)G_{0}$ when the BCS coherence peak aligns with the MZM [33, 34]. Despite the lack of topological protection in this minimal setup, we find a less obvious result: the universal conductance values through either the normal or the SC probe are robust against local detuning of the QD in contact with such probe. Further analysis in this work includes a detailed study of the role of the asymmetries and non-local effects induced by the inclusion of the non-trivial self-energy associated to the SC lead. As we discuss, this phenomenology could be used as a way to measure the degree of localization [35] of the MZMs in the minimal Kitaev chain and to characterize sweet spots.
@@
Interestingly, the regime of strong hybridization between the SC and the minimal Kitaev chain is governed by the emergence of Yu-Shiba-Rusinov (YSR) states [36, 37, 38, 39, 40, 41]. Qubits based on subgap and YSR states [42, 43, 44, 45, 46, 47, 48] have been shown to feature enhance robustness against noise [49, 50, 51]. Additionally, YSR states have been found to increase the stability of minimal Kitaev chains using toy models that neglect the SC continuum of states [52, 53]. Their inclusion, however, results in important physical consequences, since the energy of the YSR states decreases when increasing the coupling to the lead, thus reducing the effective gap between MZMs and the excited states. A finite spin-polarization in the dots generates a destructive interference between the YSR states and the MZMs, which spoils the topological properties of the latter, including the conductance quantization. In the language of non-Hermitian open quantum systems, we understand the topological properties of the hybrid junction [54, 55, 56] in terms of Exceptional Point (EP) bifurcations in the subgap spectrum [57, 58, 59, 60, 61, 62, 63], characterized by a change of the MZMs localization.
@@
We finally offer, in Sec. VII, some conclusions, summarizing the main results. Technical details like the explicit analysis of the poles of the system including the exceptional points, some notes on the Green’s functions used to compute both, spectral and transport properties and, the extension of the study of the physics associated to the QDs chemical potential to large couplings, are included in the appendices.
@@ II Modelization
We examine a minimal Kitaev chain coupled to a normal and a SC leads, as sketched in Fig.1(a). We consider that the system is subject to a strong magnetic field that fully polarizes the quantum dots spins. The two QDs are coupled via a central SC, that mediates ECT and CAR, with effective amplitudes $t$ and $\Delta_{p}$ respectively, realizing a one dimensional spinless p-wave SC that can host MZMs [12]. Written in the $4\times 4$ (hat notation) Nambu basis $\hat{\Psi}=(\psi_{\uparrow},\psi_{\uparrow}^{\dagger},\psi_{\downarrow},\psi_{\downarrow}^{\dagger})^{T}$, the toy model for the minimal Kitaev chain takes the form
$\hat{H}_{\nu}=\varepsilon_{\nu}\sigma_{z}\tau_{0}+2E_{z,\nu}\sigma_{z}\Pi_{\downarrow}\ ,\quad\hat{T}=(t\sigma_{z}+i\Delta_{p}\sigma_{y})\Pi_{\uparrow}\ ,$ (1)
where $\nu={1,2}$ corresponds to the different sites in the chain, $\tau_{\mu}$ ($\sigma_{\mu}$) are Pauli matrices acting in spin (electron/hole) space, and $\Pi_{\uparrow,\downarrow}=(\tau_{0}\pm\tau_{z})/2$ are the projectors over the spin up/down sector. Notice that as one reaches the polarized limit of the minimal Kitaev chain when the Zeeman splitting $E_{z,\nu}\rightarrow\infty$, the coupling between the QDs in the spin down sector goes correspondingly to zero, cf. App. A. External gates can be used to tune the chemical potential on the two QDs independently ($\varepsilon_{1}$ and $\varepsilon_{2}$), changing the degree of localization of the MZMs, see Fig.1(b). The minimal Kitaev chain is coupled to the SC (normal) lead through the tunneling rate $\Gamma_{s}=\lambda_{s}^{2}/t_{s}$ ($\Gamma_{n}=\lambda_{n}^{2}/t_{n}$), being $\lambda_{s}$ ($\lambda_{n}$) the coupling and $t_{s}$ ($t_{n}$) the bandwidth.
@@
To describe the transport properties of the device, we use the non-equilibrium Green’s function formalism [64, 34, 65, 66, 67], that allows us to consider arbitrary coupling strengths between the minimal Kitaev chain and the leads, see Appendices A, D and F for details. In addition, the poles of the advanced Green’s function determine the energy of the system’s states that can be related to the main transport features. Notice that for a spin-polarized chain, there are no subgap Andreev reflections with the SC probe [68] due to the incompatibility between order parameters across the junction [69, 33, 34] (i.e., a BCS lead with s-wave pairing $\Delta_{s}$ and a spinless SC with p-wave pairing $\Delta_{p}$ [70, 71]). Unless otherwise stated, we use $\Delta_{p}=\Delta_{s}=\Delta=t=1$ and $t_{s}=t_{n}=10\,t$.
@@
Figure 2: Local (a,b) and non-local (c,d) effects in SC transport in the weak coupling limit, varying the chemical potential of the dot in contact with the respective probe. First (Second) row shows the dependence of the SC conductance for various values of $\varepsilon_{1}$ ($\varepsilon_{2}$). The different columns show different coupling to the normal lead $\lambda_{n}/t=0.1$ (a,c) and $\lambda_{n}/t=0.5$ (b,d). Inset shows a diagram of the setup, where the blue arrow marks the direction of the probing. Red dashed line marks the universal height value $G_{s}/G_{0}=(4-\pi)$. We set $\lambda_{s}/t=0.5$ in all panels.
The poles of the Green’s function are contained in the characteristic polynomial $P(\omega)=g_{e}g_{h}-f^{2}$ with
$f(\omega)$ $=$ $-(\gamma_{+}+\gamma_{-})t\Delta_{p}\tau\,,$ (2)
where $\tau=\sqrt{\Delta_{s}^{2}-\omega^{2}}$ and $\gamma_{\pm}=(\omega\pm\varepsilon_{2}-i\Gamma_{n})$. In the sweet spot (i.e., $t=\Delta_{p}$ and $\varepsilon_{1}=\varepsilon_{2}=0$) and $\Gamma_{n}\rightarrow 0$ limit, the polynomial has 8th degree and simplifies to,
$P(\omega)$ $=\omega^{4}\big{(}\Gamma_{s}+\sqrt{\Delta_{s}^{2}-\omega^{2}}\big{)}\,Q(\omega)\,,$ (3a)
$Q(\omega)$ $=\omega^{2}\Gamma_{s}+(\omega^{2}-4\Delta_{p}^{2})\sqrt{\Delta_{s}^{2}-\omega^{2}}\,,$ (3b)
where the poles of $Q(\omega)$ may describe the excited states (±$\pm z_{s,e}$) in certain parameter regimes. Notice that the non-perturbative coupling to the SC lead renormalizes the energy of such excited states, appearing as subgap poles detaching from the continuum. These poles can be described analytically, and we reproduce here an approximation to the lowest contribution in $\Gamma_{s}$
$z_{s,e}\approx\frac{1}{\sqrt{3}}\sqrt{8\Delta_{p}^{2}+\Delta_{s}^{2}-\alpha^{1/3}(\Gamma_{s})-\frac{\delta^{2}}{\alpha^{1/3}(\Gamma_{s})}}\;,$ (4)
where $\alpha(\Gamma_{s})=\big{(}\delta^{3}-12\sqrt{3}\,\Gamma_{s}\Delta_{p}^{2}\,\delta^{3/2}\big{)}$, and $\delta=(4\Delta_{p}^{2}-\Delta_{s}^{2})$, cf. App. A for details.
@@ III Weak coupling
Figure 4: (a) QD2 LDOS as a function of the hybridization parameter to the SC lead. (b,e) LDOS across the junction, showing the spreading of the states into the SC lead in the $\lambda_{n}/t\rightarrow 0$ limit. (c,f) LDOS cuts following the YSR state and the MZM respectively. (b,c) [(e,f)] Strong (Weak) coupling for $\lambda_{s}/t=4$ ($\lambda_{s}/t=0.5$) marked by arrows in (a). (d, g) Illustration of the weak and strong coupling across the junction (d) Weak coupling, minimal Kitaev chain. (g) Strong coupling, YSR state coupled to QD2 forming bonding and anti-bonding subgap states. White dashed line in panel (a) represents the analytic pole obtained from Eq. (II) when considering $\Gamma_{n}\neq 0$. The LDOS at QD1 and QD2 in (c,f) is rescaled for visualization purposes. We set $\lambda_{n}/t=1$ in the rest of the panels.
Figure 1(b) illustrates the Majorana localization when detuning the two QDs away from the sweet spot. When QD1 is detuned, the two MZMs have a finite spectral weight in QD2. Thus, when considering $\varepsilon_{1}\neq 0$, the normal lead couples to both MZMs, producing a destructive interference sharp dip at zero energy, Fig.1(c), as previously observed in Ref. [12]. Moreover, due to electrons tunneling between the normal and the SC lead, mediated by the Kitaev chain, a peak appears at $e|V|=\Delta$, where $\Delta_{p}=\Delta_{s}=\Delta$, see also Fig.1(d).
By contrast, when QD2 is detuned from the sweet spot, the two MZMs acquire a finite weight in QD1. This leads to a quantized zero-bias conductance peak, showing a width that depends on $\varepsilon_{2}$, cf. Fig.1(f) and Ref. [12]. Therefore, detuning either the QD attached to the normal lead or the one attached to the SC lead has very different effects on the transport properties.
@@
Figure 2(c,d) illustrates the opposite behaviour, namely how SC transport is modified non-locally when the QD far from the SC lead is detuned. In this case, the lead couples to both MZMs. The resulting hybridization becomes evident when testing transport through the SC probe, where the quantized conductance value gets reduced as soon as $\varepsilon_{2}\neq 0$. Moreover, the coherent coupling between MZMs mediated by the SC lead, gives rise to subgap signatures in transport, Fig. 2(c). We remark that the subgap structure is not associated to any renormalization of subgap poles but to inelastic single-particle processes in transport, cf. Refs. [72] and App. D. When the coupling to the normal lead is increased, the spreading of the MZM into the latter reduces the hybridization between MZMs and therefore, enhances the SC transport, Fig. 2(d). Notice that this dependence with $\Gamma_{n}$ is the opposite to the one found in the $\varepsilon_{1}\neq 0$ case in Fig. 2(a, b).
@@ IV Strong coupling
Figure 5: (a) QD2 LDOS as a function of the hybridization parameter to the SC lead increasing from weak to the ultra-strong coupling limit. First row, LDOS across the junction, showing the spreading of the states into the SC lead. Second row, real and imaginary part of the poles of the system (solid and dotted lines respectively) for $\lambda_{s}/t=10,60,100$ marked by arrows in (a). White dashed line in panel (a) represents the analytic pole obtained from Eq. (II) when considering $\Gamma_{n}\neq 0$. We set $\lambda_{n}/t=1$ in all panels, except indicated, we have set $t/\Delta=1$.
Now we examine the electron transport as a function of the coupling to the SC lead, going beyond the weak coupling regime ($\Gamma_{s}/t\gtrsim 1$) where the excited state $z_{s,e}=2\Delta$ is strongly renormalized by the lead, emerging as a subgap YSR state [72, 73, 74, 75, 52, 53]. Transport provides an ideal method to observe the interplay between the aforementioned YSR states and MZMs, Fig. 3. The onset of the YSR state appears as a new peak detaching from the continuum with linear dispersion around $\Gamma_{s}/t\gtrsim 0.5$, Figs. 3(a, c). This phenomenology does not affect the MZM signatures at $V\ll\Delta$, except for $\Gamma_{s}/t\gg 1$ and finite spin-polarization in QD1, see below and App. C. The conductance of the excited state is also quantized for $\Gamma_{s}/t\geq 1$ and $\varepsilon_{2}=0$. Intuitively, the excited state is no longer broadened by the SC lead when is well detached from the continuum, thus recovering the Majorana properties of the original excited state.
Detuning $\varepsilon_{1}$ barely affects the excited state and only leads to a dip in the zero-bias conductance, as commented in the weak coupling regime, see Fig. 1. The energy of the excited state is renormalized when $\varepsilon_{2}$ shifts away from the sweet spot, losing any quantized behaviour in transport. In particular, it gets pinned to $z_{s,e}\rightarrow\varepsilon_{2}$ for large $\Gamma_{s}$ values, see Fig. S 2 in App. A.
@@
Transport through the SC lead is also affected when $\Gamma_{s}$ increases, Fig. 3(b, d). First, the universal height value $(4-\pi)$ moves to larger values of the voltage bias as the coupling to the SC lead is increased [72], illustrated by the yellow line. Figure 3(d) shows the onset of a subgap structure associated to the hybridization between Majoranas when detuning QD2 even in the weak coupling regime $\Gamma_{s}\rightarrow 0$, see also Fig. 2(c). This subgap structure is obtained from the single-particle contribution when testing transport through the SC probe when enabling thermally activated inelastic cotunneling. Thus, the hybridized MZMs behaves as a trivial zero energy subgap state in the strong tunneling regime [72], cf. App. D and E. Notice that the conductance peaks have a reduced value, compared to panel (b) and for $\Gamma_{s}/t\leq 1$ the dispersion of the subgap peak is linear with $\lambda_{s}$ (i.e., quadratic with $\Gamma_{s}$) highlighting the different origin with the observed subgap structure in Fig. 3(a,c).